# Learning from Linear Algebra: A Graph Neural Network Approach to Preconditioner Design for

### v2

@@ Abstract
Large linear systems are ubiquitous in modern computational science and engineering. The main recipe for solving them is the use of Krylov subspace iterative methods with well-designed preconditioners. Deep learning models can be used as nonlinear preconditioners during the iteration of linear solvers such as the conjugate gradient (CG) method. Neural network models require an enormous number of parameters to approximate well in this setup. Another approach is to take advantage of small graph neural networks (GNNs) to construct preconditioners with predefined sparsity patterns. Recently, GNNs have been shown to be a promising tool for designing preconditioners to reduce the overall computational cost of iterative methods by constructing them more efficiently than with classical linear algebra techniques. However, preconditioners designed with these approaches cannot outperform those designed with classical methods in terms of the number of iterations in CG. In our work, we recall well-established preconditioners from linear algebra and use them as a starting point for training the GNN to obtain preconditioners that reduce the condition number of the system more significantly. Numerical experiments show that our approach outperforms both classical and neural network-based methods for an important class of parametric partial differential equations. We also provide a heuristic justification for the loss function used and show that preconditioners obtained by learning with this loss function reduce the condition number in a more desirable way for CG.
@@ 1 Introduction
Typically, the application of parametric PDEs produces large linear systems, often with entries of varying scale, and therefore poses significant computational challenges. Projection Krylov subspace iterative methods are widely used to solve such systems. They rely on finding an optimal solution in a subspace constructed as follows: $\mathcal{K}_{r}(A,b)=\text{span}\{b,~{}Ab,~{}A^{2}b,\dots,~{}A^{r-1}\}.$.
The conjugate gradient (CG) method is used to solve large sparse systems with symmetric and positive definite (SPD) matrices (Saad, 2003; Axelsson, 1996). CG has a well-established convergence theory and convergence guarantees for any symmetric matrix. However, the convergence rate of CG is determined by $\sqrt{\kappa(A)}$. The condition number $\kappa(A)$ of an SPD matrix $A$, defined in $2\text{-norm}$, is a ratio between the maximum and minimum eigenvalues $\kappa(A)=|\lambda_{\max}|\big{/}|\lambda_{\min}|$.
Real-world applications with non-smooth high-contrast coefficient functions and high-dimensional linear systems separate eigenvalues and results into ill-conditioned problems. Decades of research in numerical linear algebra have been devoted to constructing preconditioners $P$ for ill-conditioned $A$ to improve the condition number in the form (for left-preconditioned systems) $\kappa(P^{-1}A)\ll\kappa(A)$.
@@
In this paper, we propose a novel neural method for preconditioner design called PreCorrector (short for Preconditioner Corrector). Preconditioners constructed with PreCorrector have better effect on the spectrum than classical preconditioners by learning correction for the latter. Our contributions are as follows:
@@
    We propose a novel scheme for preconditioner design based on learning correction for well-established preconditioners from linear algebra with the GNN.
@@ 2 Neural design of preconditioner
In this paper we focus on the SPD matrices so instead of ILU, ILU($p$) and ILUt($p$) we use the incomplete Choletsky factorization IC, IC($p$) and ICt($p$). Further, we will form the preconditioners in the form of Choletsky decomposition (Trefethen & Bau, 2022) $P=LL^{\top}$ with sparse $L$ obtained by different methods.
@@
Our utlimate goal is to find such a decomposition that $\kappa((L(\theta)L(\theta)^{\top})^{-1}A)\ll\kappa((LL^{\top})^{-1}A)\ll\kappa(A)$, where $L$ is the classical numerical IC decomposition and $L(\theta)=\mathcal{F}(A)$ is an approximate decomposition with some function $\mathcal{F}$. Several papers (Li et al., 2023; Häusner et al., 2023) suggest using GNN as a function $\mathcal{F}$ to minimize certain loss function:
@@
$\big{\|}\big{(}P-A\big{)}A^{-1}\big{\|}^{2}_{F}=\big{\|}PA^{-1}-I\big{\|}_{F}^{2}=\operatorname{Tr}\Big{(}(PA^{-1}-I)^{\top}(PA^{-1}-I)\Big{)}\\ =\mathbb{E}_{\varepsilon}\Big{[}\varepsilon^{\top}(PA^{-1}-I)^{\top}(PA^{-1}-I)\varepsilon\Big{]}=\mathbb{E}_{\varepsilon}\big{\|}(PA^{-1}-I)\varepsilon\big{\|}_{2}^{2},\quad\varepsilon\sim\mathcal{N}(0,1).$ (4)
@@
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\big{\|}L(\theta)L(\theta)^{\top}x_{i}-b_{i}\big{\|}_{2}^{2}$ (5)
@@ 3 Learn correction for ILU
Our main goal is to construct preconditioners that reduce the condition number of an SPD matrix more than classical preconditioners with the same sparisy pattern.
3.1 Graph neural network with preserving sparsity pattern
Following the idea of (Li et al., 2023), we use the message-passing GNN architecture (Zhou et al., 2020) to preserve the sparsity pattern and predict the lower triangular matrix to create a preconditioner in a form of IC decomposition.
@@ 3.1 Graph neural network with preserving sparsity pattern
The duality between sparse matrices and graphs is used to obtain vertices and edges, such as $Ax=b\rightarrow\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $a_{i,j}=e_{i,j}\in\mathcal{E},b_{i}=v_{i}\in\mathcal{V}$. The original GNN architecture from (Li et al., 2023):
  • 1.
    First step is to use node and edge encoders to increase their dimensionality with multi-layer perceptrons (MLPs): $v_{i}=\text{MLP}_{v}(v_{i}),e_{i,j}=\text{MLP}_{e}(e_{i,j})$.
  • 2.
    Then the encoded graph is processed with $T$ rounds of message passing (Brandstetter et al., 2022) ($t=1,\dots,T$) to transfer information between vertices and edges. During a single round, we update vertices with $v_{i,t+1}=\text{MLP}_{mp,v}(v_{i,t},\sum_{j}e_{i,j,t}v_{j,t})$, and then update the edges with $e_{i,j,t+1}=\text{MLP}_{mp,e}(e_{i,j,t},v_{i,t+1}v_{j,t+1}),\,\,\text{for}\,\,i\neq j$.
  • 3.
    Next step is to decode the lower triangular matrix while preserving the information in the upper triangular part of the matrix. To do this we average the bidirectional edges, decode them with MLP and then zero out the upper triangular part: $e_{i,j,T}=\big{(}e_{i,j,T}+e_{j,i,T}\big{)}/2$ and $L_{i,j|i\leq j}=\text{MLP}_{\text{decod}}(e_{i,j,T}),L_{i,j|i>j}=0$.
  • 4.
    After all round of message passing the diagonal of the decomposition inherited as the diagonal from original matrix to ensure SPD property in resulting decomposition $\text{diag}(L(\theta)):=\sqrt{\text{diag}(A)}$.
  • 5.
    Finally, assemble the preconditioner in a form of Choletsky decomposition $P:=L(\theta)L(\theta)^{\top}$.
In our experiments, we observe that training GNN from scratch can be unstable and results in preconditioners that have weaker effect on the spectrum than their classical analogues. Moreover, with growing matrix size, the very first step of training, when GNN is initialized with random weights, overflows loss since the residual with random $P$ is huge. We propose to solve both problems by learning corrections to classical preconditioners.
3.2 PreCorrector
@@
Instead of passing left hand side matrix $A$ as input to GNN in (1), we propose: (i) to pass $L$ from the IC decomposition to the GNN and (ii) to train GNN to predict a correction for this decomposition (Figure 1):
$L(\theta)=L+\alpha\cdot\text{GNN}(\theta,L,b).$ (6)
@@
The correction coefficient $\alpha$ is also a learning parameter that is updated during training. At the beginning of training, we set $\alpha=0$ to ensure that the first gradient updates come from pure IC factorization. Since we already start with a good initial guess, we observed that pinning the diagonal is redundant and limits the training of the PreCorrector. Moreover, GNN in (6) takes as input the lower-triangular matrix $L$ from IC instead of $A$, so we are not anchored to a single specific sparse pattern of $A$ and we can: (i) omit half of the graph and speed up the training process and (ii) use different sparsity patterns to obtain even better preconditioners. In Experiment section, we show that the proposed approach with input $L$ from IC(0) and ICt(1) produces better preconditioners compared to classical IC(00) and ICt($1$) and previous preconditioners using neural networks.
[…]
Figure 1: PreCorrector scheme that takes IC(00) as input. (a) Start with linear system $Ax=b$. (b) Obtain $L_{0}$ from IC(0) decomposition. (c) Input $L_{0}$ and $b$ to GNN. (d) Calculate $L(\theta)$ with equation 6 and construct preconditioner in form of IC. (Very right picture) Note that obtained preconditioner in form of IC(0) decomposition can be stored as initial matrix $A$.
@@ 4 Dataset
The diffusion equation is chosen because it occurs frequently in many engineering applications, such as: composite modeling (Carr & Turner, 2016), geophysical surveys (Oristaglio & Hohmann, 1984), fluid flow modeling (Muravleva et al., 2021). In these industrial applications, the coefficient functions are discontinuous, i.e. they change rapidly within neighbouring cells. Examples include the flow of immiscible fluids of different viscosities and fluid flow in heterogeneous porous media.
@@
We propose to use a Gaussian Random Field (GRF) $\phi(x)$ to generate the coefficients. To control the complexity of the diffusion equation with discontinuous coefficients, we can measure the contrast in the GRF:
@@
Then we generate coefficients for the equation (7) as $k(x)=\exp{(\phi(x))}$.
By changing the variance in the GRF, we can control the contrast of the coefficients and thus the complexity. We generate datasets for each grid value from $\{32,~{}64,~{}128\}$. The contrast in the diffusion equation is controlled with a variance in the coefficient function GRF and takes value from $\{0.1,~{}0.5,~{}0.7\}$. The Gaussian covariance model is used in GRF. The forcing term $f$ is sampled from the standard normal distribution $\mathcal{N}(0,~{}1)$ and each PDE is discretized using the 5-point finite difference method. GRF is generated using the parafields library [^1] https://github.com/parafields/parafields. More details about datasets can be found in the Appendix A.6.
@@ 5 Experiments
The main idea behind using GNNs to construct preconditioners is to preserve the sparsity pattern. Therefore, the algorithmic complexity of using preconditioners (matrix-vector product) is the same when using preconditioners with the same sparsity pattern. This allows a fair evaluation of the quality of neural preconditioners with the same sparsity pattern only in terms of the number of CG iterations. Furthermore, all approaches to IC decomposition with the same sparsity pattern, including the classical ones, compete with each other in terms of construction time, effect on the spectrum (i.e., number of CG iterations) and generalization ability.
@@
Table 1: Comparison on diffusion equation with variance $0.7$: classical algorithms for IC(00) and ICt($1$) and PreCorrector. Pre-time stands for precomputations time.
Time (iters) Time (iters) Time (iters)
Grid Method Pre-time to $10^{-3}$ to $10^{-6}$ to $10^{-9}$
$64\times 64$ $\text{IC}(0)$ $1.6\cdot 10^{-4}$ $0.169\pm 0.006$ $0.213\pm 0.007$ $0.248\pm 0.007$
($77\pm 2.9$) ($98\pm 3.2$) ($115\pm 3.1$)
PreCor$\big{[}\text{IC}(0)\big{]}$ $2.0\cdot 10^{-3}$ $\textbf{0.112}\pm 0.097$ $\textbf{0.141}\pm 0.098$ $\textbf{0.170}\pm 0.099$
($42\pm 1.7$) ($55\pm 2.1$) ($67\pm 2.5$)
$\text{ICt}(1)$ $8.7\cdot 10^{-4}$ $0.105\pm 0.004$ $0.133\pm 0.004$ $0.155\pm 0.004$
($46\pm 1.8$) ($59\pm 1.9$) ($69\pm 1.9$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $2.3\cdot 10^{-3}$ $\textbf{0.091}\pm 0.070$ $\textbf{0.115}\pm 0.071$ $\textbf{0.138}\pm 0.072$
($29\pm 1.4$) ($38\pm 1.7$) ($47\pm 1.9$)
$128\times 128$ $\text{IC}(0)$ $5.1\cdot 10^{-4}$ $1.071\pm 0.189$ $1.338\pm 0.229$ $1.554\pm 0.261$
($156\pm 5.6$) ($196\pm 5.5$) ($228\pm 5.7$)
PreCor$\big{[}\text{IC}(0)\big{]}$ $2.0\cdot 10^{-3}$ $\textbf{0.571}\pm 0.078$ $\textbf{0.720}\pm 0.079$ $\textbf{0.859}\pm 0.079$
($67\pm 3.2$) ($85\pm 3.5$) ($102\pm 3.9$)
$\text{ICt}(1)$ $3.8\cdot 10^{-3}$ $0.720\pm 0.098$ $0.902\pm 0.122$ $1.048\pm 0.141$
($95\pm 3.5$) ($119\pm 3.4$) ($139\pm 3.5$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $5.4\cdot 10^{-3}$ $\textbf{0.470}\pm 0.071$ $\textbf{0.593}\pm 0.078$ $\textbf{0.708}\pm 0.086$
($50\pm 2.5$) ($64\pm 2.9$) ($77\pm 3.1$)
Table 2: Comparison on diffusion equation with variance $0.7$: classical algorithms for ICt($5$) and PreCor$\big{[}\text{ICt}(1)\big{]}$. Pre-time stands for precomputations time.
Time (iters) Time (iters) Time (iters)
Grid Method Pre-time to $10^{-3}$ to $10^{-6}$ to $10^{-9}$
$64\times 64$ $\text{ICt}(5)$ $2.1\cdot 10^{-3}$ $\textbf{0.069}\pm 0.003$ $\textbf{0.087}\pm 0.003$ $\textbf{0.103}\pm 0.003$
($22\pm 0.8$) ($28\pm 0.8$) ($34\pm 0.8$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $2.3\cdot 10^{-3}$ $0.091\pm 0.070$ $0.115\pm 0.071$ $0.138\pm 0.072$
($29\pm 1.4$) ($38\pm 1.7$) ($47\pm 1.9$)
$128\times 128$ $\text{ICt}(5)$ $8.9\cdot 10^{-3}$ $0.570\pm 0.071$ $0.705\pm 0.088$ $0.833\pm 0.104$
($48\pm 1.7$) ($60\pm 1.6$) ($70\pm 1.8$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $5.4\cdot 10^{-3}$ $\textbf{0.470}\pm 0.071$ $\textbf{0.593}\pm 0.078$ $\textbf{0.708}\pm 0.086$
($50\pm 2.5$) ($64\pm 2.9$) ($77\pm 3.1$)
Experiments environment
Each dataset in the Section 4 consists of $1000$ training and $200$ test linear systems. The final neural networks are trained with batch size $8$, learning rate $10^{-3}$ and Adam optimizer. For a fair comparison, we set the GNN architecture to $5$ message passing rounds and $2$ hidden layers with $16$ hidden features in all MLPs (see Section 3.1) in each experiment. PreCorrector training always starts with the parameter $\alpha=0$ in (6). For GNN training, we used libraries from the JAX ecosystem: jax (Bradbury et al., 2018), optax (Kidger & Garcia, 2021), equinox (DeepMind et al., 2020). We used a single GPU Nvidia A40 48Gb for training. The construction time of preconditioners with neural design was measured on the same GPU. Preconditioners with classical algorithms were generated on the Intel(R) Xeon(R) Gold 6342 CPU @ 2.80GHz with ilupp library [^2] https://github.com/c-f-h/ilupp Mayer (2007). The CG method was run on the same CPU using the scipy (Virtanen et al., 2020) implementation.

### v3

@@ Abstract
Large linear systems are ubiquitous in modern computational science and engineering. The main recipe for solving them is the use of Krylov subspace iterative methods with well-designed preconditioners. Recently, GNNs have been shown to be a promising tool for designing preconditioners to reduce the overall computational cost of iterative methods by constructing them more efficiently than with classical linear algebra techniques. Preconditioners designed with these approaches cannot outperform those designed with classical methods in terms of the number of iterations in CG. In our work, we recall well-established preconditioners from linear algebra and use them as a starting point for training the GNN to obtain preconditioners that reduce the condition number of the system more significantly than classical preconditioners. Numerical experiments show that our approach outperforms both classical and neural network-based methods for an important class of parametric partial differential equations. We also provide a heuristic justification for the loss function used and show that preconditioners obtained by learning with this loss function reduce the condition number in a more desirable way for CG.
Machine Learning, ICML
@@ 1 Introduction
Typically, the application of parametric PDEs produces large linear systems, often with entries of varying scale, and therefore poses significant computational challenges. Projection Krylov subspace iterative methods are widely used to solve such systems. They rely on finding an optimal solution in a subspace constructed as follows: $\mathcal{K}_{r}(A,b)=\text{span}\{b,~{}Ab,~{}A^{2}b,\dots,~{}A^{r-1}\}$.
The conjugate gradient (CG) method is used to solve large sparse systems with symmetric and positive definite (SPD) matrices (Saad, 2003; Axelsson, 1996). CG has a well-established convergence theory and convergence guarantees for any symmetric matrix. However, the convergence rate of CG is determined by $\sqrt{\kappa(A)}$, which is typically large. The condition number $\kappa(A)$ of an SPD matrix $A$, defined in $2\text{-norm}$, is a ratio between the maximum and minimum eigenvalues $\kappa(A)=|\lambda_{\max}|\big{/}|\lambda_{\min}|$.
Real-world applications with non-smooth high-contrast coefficient functions and high-dimensional linear systems separate eigenvalues and results into ill-conditioned problems with high $\kappa(A)$. Decades of research in numerical linear algebra have been devoted to constructing preconditioners $P$ for ill-conditioned $A$ to improve the condition number in the form (for left-preconditioned systems) $\kappa(P^{-1}A)\ll\kappa(A)$.
@@
In this paper, we propose a novel neural method of preconditioner design for symmetric positive definite matrices called PreCorrector (short for Preconditioner Corrector). Preconditioners constructed with PreCorrector have better effect on the spectrum than classical preconditioners by learning correction for the latter. Our contributions are as follows:
@@
    We propose a novel scheme for preconditioner design based on learning correction for well-established preconditioners from linear algebra with the GNN, which has better effect on spectrum than classical preconditioners.
@@ 2 Neural design of preconditioner
In this paper we focus on the SPD matrices so instead of ILU, ILU($p$) and ILUt($p$) we use the incomplete Cholesky factorization IC, IC($p$) and ICt($p$). Further, we will form the preconditioners in the form of Cholesky decomposition (Trefethen & Bau, 2022) $P=LL^{\top}$ with sparse $L$ obtained by different methods.
@@
Our ultimate goal is to find such a decomposition that $\kappa((L(\theta)L(\theta)^{\top})^{-1}A)\ll\kappa((LL^{\top})^{-1}A)\ll\kappa(A)$, where $L$ is the classical numerical IC decomposition and $L(\theta)=\mathcal{F}(A)$ is an approximate decomposition with some function $\mathcal{F}$. Several papers (Li et al., 2023; Häusner et al., 2023) suggest using GNN as a function $\mathcal{F}$ to minimize certain loss function:
@@
$\big{\|}\big{(}P-A\big{)}A^{-1}\big{\|}^{2}_{F}=\big{\|}PA^{-1}-I\big{\|}_{F}^{2}=\\ \operatorname{Tr}\Big{(}(PA^{-1}-I)^{\top}(PA^{-1}-I)\Big{)}=\\ \mathbb{E}_{\varepsilon}\Big{[}\varepsilon^{\top}(PA^{-1}-I)^{\top}(PA^{-1}-I)\varepsilon\Big{]}=\\ \mathbb{E}_{\varepsilon}\big{\|}(PA^{-1}-I)\varepsilon\big{\|}_{2}^{2},\quad\varepsilon\sim\mathcal{N}(0,1).$ (4)
@@
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\big{\|}L(\theta)L(\theta)^{\top}x_{i}-b_{i}\big{\|}_{2}^{2}~{}.$ (5)
@@ 3 Learn correction for ILU
Our main goal is to construct preconditioners that reduce the condition number of an SPD matrix more than classical preconditioners with the same sparsity pattern. Several previous papers suggested to use convolutional neural networks with sparse convolutions (Sappl et al., 2019; Calì et al., 2023) and graph neural networks (Li et al., 2023; Häusner et al., 2023). However, none of the approaches, to the best of authors knowledge, did not achieve to obtain better effect of spectrum with neural-preconditioner construction while maintaining the same sparsity level. We propose to learn corrections for classical preconditioner to achieve such effect.
Figure 1: Inplace updating of IC(00) factor allows to obtain better preconditioner.
Inplace ILU update
PreCorrector is built around the idea that classical preconditioners of the ILU family produce reliable preconditioners for the CG method. Although the algorithm for their construction is known, they do not exist uniquely and one can try to train neural network to construct preconditioners of the same sparsity pattern with better effect on spectrum. This can be shown by inplace updating of IC(0) factors with gradient descent (Figure 1).
In this approach we treat each element of the sparse matrix as a neural network parameter and update an already efficient preconditioner. We follow the idea of recent research and use GNN to parametrize preconditioner learning.
PreCorrector
@@
The duality between sparse matrices and graphs is used to obtain vertices and edges, such as $Ax=b\rightarrow\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $a_{i,j}=e_{i,j}\in\mathcal{E},b_{i}=v_{i}\in\mathcal{V}$.
@@
Instead of passing left hand side matrix $A$ as input to GNN in (1), we propose: (i) to pass $L$ from the IC decomposition to the GNN and (ii) to train GNN to predict a correction for this decomposition:
$L(\theta)=L+\alpha\cdot\text{GNN}(\theta,L)~{}.$ (6)
@@
The correction coefficient $\alpha$ is also a learning parameter that is updated during training. At the beginning of training, we set $\alpha=0$ to ensure that the first gradient updates come from pure IC factorization. Moreover, GNN in (6) takes as input the lower-triangular matrix $L$ from IC instead of $A$, so we are not anchored to a single specific sparse pattern of $A$ and we can: (i) omit half of the graph and speed up the training process and (ii) use different sparsity patterns to obtain even better preconditioners. In Experiment section, we show that the proposed approach with input $L$ from IC(0) and ICt(1) produces better preconditioners compared to classical IC(00) and ICt($1$) and previous neural-designed preconditioners.
Graph neural network architecture
We use ubiquitous encode-process-decode (Battaglia et al., 2018) architecture for the GNN, where multilayer perceptrons are used for encoders and decoders. Process-block consists of message-passing layer that increase receptive field of the GNN with multiple rounds. Before encoder and after decoder we apply normalization and renormalization for the edges respectively. One can see PreCorrector’s forward pass in Algorithm 1.
Note that in (6) we omit the right-hand side vector $b$ and do not use it as input for GNN. In our experiments, we observe that GNN with the right-hand side as node input does not effect resulting preconditioner quality, while it suffers from sensitivity to the right-hand side vector. Although we can either use a different representation of the node features (e.g., as in (Häusner et al., 2023)) or use diagonal elements of the matrix. Instead we set vector $\textbf{1}=[1,\dots,1]^{T}$ as the node input to the message passing layer, which allowed us to reduce the number of parameters by a factor of $2.3$ while maintaining the quality of the resulting preconditioner. Details on the architecture of the PreCorrector architecture can be found in the Appendix A.2.
@@ 4 Dataset
The diffusion equation is chosen because it occurs frequently in many engineering applications, such as: composite modeling (Carr & Turner, 2016), geophysical surveys (Oristaglio & Hohmann, 1984), fluid flow modeling (Muravleva et al., 2021). In these industrial applications, the coefficient functions are discontinuous, i.e. they change rapidly within neighbouring cells. Examples also include the flow of immiscible fluids of different viscosities and fluid flow in heterogeneous porous media.
@@
We propose to use a Gaussian Random Field (GRF) $\phi(x)$ to generate the coefficients. In our experiments we use the parafields library [^1] https://github.com/parafields/parafields to generate GRFs. To control the complexity of the diffusion equation with discontinuous coefficients, we can measure the contrast in the GRF:
@@
Then we generate coefficients for the equation (7) as $k(x)=\exp{(\phi(x))}$. By changing the variance in the GRF, we can control the contrast of the coefficients and thus the complexity. More details about datasets can be found in the Appendix A.3.
Table 1: Comparison on diffusion equation with variance $1.1$: classical algorithms IC(00) and ICt($1$) and PreCorrector. Lower is better, the best results are bold. Pre-time stands for precomputations time.
Time (iters) Time (iters) Time (iters) Time (iters)
Grid Method Pre-time to $10^{-3}$ to $10^{-6}$ to $10^{-9}$ to $10^{-12}$
$128\times 128$ $\text{IC}(0)$ $4.6\cdot 10^{-4}\pm 9.4\cdot 10^{-6}$ $1.267\pm 0.154$ $1.573\pm 0.186$ $1.816\pm 0.218$ $2.018\pm 0.243$
($171\pm 6.8$) ($213\pm 7.3$) ($245\pm 7.0$) ($273\pm 6.8$)
PreCor$\big{[}\text{IC}(0)\big{]}$ $8.4\cdot 10^{-4}\pm 1.3\cdot 10^{-4}$ $\textbf{0.763}\pm 0.142$ $\textbf{0.969}\pm 0.180$ $\textbf{1.164}\pm 0.218$ $\textbf{1.355}\pm 0.254$
($108\pm 10.3$) ($138\pm 13.4$) ($166\pm 16.5$) ($193\pm 19.6$)
$\text{ICt}(1)$ $3.1\cdot 10^{-3}\pm 6.2\cdot 10^{-5}$ $0.839\pm 0.121$ $1.040\pm 0.148$ $1.201\pm 0.171$ $1.337\pm 0.190$
($104\pm 4.1$) ($129\pm 4.5$) ($150.0\pm 4.2$) ($166\pm 4.0$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $3.6\cdot 10^{-3}\pm 2.3\cdot 10^{-4}$ $\textbf{0.450}\pm 0.039$ $\textbf{0.572}\pm 0.042$ $\textbf{0.687}\pm 0.049$ $\textbf{0.798}\pm 0.054$
($63\pm 3.8$) ($80\pm 4.5$) ($96\pm 5.2$) ($112\pm 5.8$)
$256\times 256$ $\text{IC}(0)$ $1.8\cdot 10^{-3}\pm 6.0\cdot 10^{-5}$ $8.638\pm 2.000$ $10.635\pm 2.355$ $12.249\pm 2.654$ $13.608\pm 2.905$
($345\pm 13.5$) ($425\pm 14.9$) ($490\pm 14.9$) ($544\pm 14.4$)
PreCor$\big{[}\text{IC}(0)\big{]}$ $3.3\cdot 10^{-3}\pm 6.1\cdot 10^{-4}$ $\textbf{3.388}\pm 0.544$ $\textbf{4.199}\pm 0.663$ $\textbf{4.943}\pm 0.775$ $\textbf{5.637}\pm 0.875$
($144\pm 12.6$) ($179\pm 15.5$) ($211\pm 17.7$) ($241\pm 19.6$)
$\text{ICt}(1)$ $1.2\cdot 10^{-2}\pm 1.8\cdot 10^{-4}$ $5.278\pm 0.593$ $6.515\pm 0.709$ $7.511\pm 0.809$ $8.340\pm 0.890$
($209\pm 8.4$) ($259\pm 9.0$) ($298\pm 9.1$) ($332\pm 8.8$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $1.4\cdot 10^{-2}\pm 6.8\cdot 10^{-4}$ $\textbf{2.719}\pm 0.381$ $\textbf{3.378}\pm 0.462$ $\textbf{3.976}\pm 0.539$ $\textbf{4.534}\pm 0.608$
($105\pm 8.3$) ($131\pm 9.8$) ($154\pm 11.2$) ($176\pm 12.4$)
Table 2: Comparison on diffusion equation with variance $1.1$: classical algorithm ICt($5$) and PreCor$\big{[}\text{ICt}(1)\big{]}$. Lower is better, the best results are bold. Pre-time stands for precomputations time.
Time (iters) Time (iters) Time (iters) Time (iters)
Grid Method Pre-time to $10^{-3}$ to $10^{-6}$ to $10^{-9}$ to $10^{-12}$
$128\times 128$ $\text{ICt}(5)$ $7.5\cdot 10^{-3}\pm 3.4\cdot 10^{-4}$ $0.716\pm 0.054$ $0.890\pm 0.062$ $1.033\pm 0.070$ $1.152\pm 0.075$
($52\pm 2.1$) ($65\pm 2.1$) ($76\pm 2.0$) ($84\pm 2.0$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $3.6\cdot 10^{-3}\pm 2.3\cdot 10^{-4}$ $\textbf{0.450}\pm 0.034$ $\textbf{0.572}\pm 0.042$ $\textbf{0.687}\pm 0.049$ $\textbf{0.798}\pm 0.054$
($63\pm 3.8$) ($80\pm 4.5$) ($96\pm 5.2$) ($112\pm 5.8$)
$256\times 256$ $\text{ICt}(5)$ $3.0\cdot 10^{-2}\pm 7.5\cdot 10^{-4}$ $5.031\pm 0.622$ $6.222\pm 0.754$ $7.178\pm 0.852$ $7.987\pm 0.942$
($110\pm 4.4$) ($136\pm 4.6$) ($157\pm 4.6$) ($175\pm 4.6$)
PreCor$\big{[}\text{ICt}(1)\big{]}$ $1.4\cdot 10^{-2}\pm 6.8\cdot 10^{-4}$ $\textbf{2.719}\pm 0.381$ $\textbf{3.378}\pm 0.462$ $\textbf{3.976}\pm 0.539$ $\textbf{4.534}\pm 0.608$
($105\pm 8.3$) ($131\pm 9.8$) ($154\pm 11.2$) ($176\pm 12.4$)
@@ 5 Experiments
The GNNs is chosen to construct preconditioners because it allows preserve the sparsity pattern. Therefore, the algorithmic complexity of using preconditioners (matrix-vector product) is the same when using preconditioners with the same sparsity pattern. This allows a fair evaluation of the quality of neural preconditioners with the same sparsity pattern only in terms of the number of CG iterations. Furthermore, all approaches to IC decomposition with the same sparsity pattern, including the classical ones, compete with each other in terms of construction time, effect on the spectrum (i.e., number of CG iterations) and generalization ability.
Table 3: Comparison on Poisson equation: classical algorithm IC(00), GNN from (Li et al., 2023) and PreCorrector. Lower is better, the best results are bold. The second best is underlined. Pre-time stands for precomputations time.
Time (iters) Tme (iters) Time (iters) Time (iters)
Grid Method Pre-time to $10^{-3}$ to $10^{-6}$ to $10^{-9}$ to $10^{-12}$
$128\times 128$ $\text{IC}(0)$ $4.7\cdot 10^{-4}\pm 2.0\cdot 10^{-5}$ $0.762\pm 0.116$ $1.012\pm 0.153$ $1.154\pm 0.175$ $1.353\pm 0.205$
($118\pm 0.0$) ($157\pm 0.0$) ($179\pm 0.0$) ($210\pm 0.0$)
PreCor$\big{[}\text{IC}(0)\big{]}$ $8.1\cdot 10^{-4}\pm 1.2\cdot 10^{-4}$ $\textbf{0.403}\pm 0.009$ $\textbf{0.520}\pm 0.009$ $\textbf{0.620}\pm 0.010$ $\textbf{0.729}\pm 0.011$
($48\pm 0.0$) ($62\pm 0.0$) ($74\pm 0.0$) ($87\pm 0.0$)
(Li et al., 2023) $1.1\cdot 10^{-3}\pm 2.5\cdot 10^{-4}$ $\underline{0.471}\pm 0.058$ $\underline{0.589}\pm 0.069$ $\underline{0.708}\pm 0.081$ $\underline{0.824}\pm 0.089$
($51\pm 0.0$) ($64\pm 0.0$) ($77\pm 0.0$) ($90\pm 0.3$)
@@
Table 4: Comparison on diffusion equation with grid $32$ with varying variance in GRF: classical algorithms for IC(00) and GNN from (Li et al., 2023). Lower is better, the best results are bold. Pre-time stands for precomputations time. 1)1)1 )None of the test linear systems converged to $10^{-3}$ tolerance in $300$ iterations.
Time (iters) Time (iters) Time (iters) Time (iters)
Variance Method Pre-time to $10^{-3}$ to $10^{-6}$ to $10^{-9}$ to $10^{-12}$
$0.1$ $\text{IC}(0)$ $7.3\cdot 10^{-5}\pm 2.7\cdot 10^{-6}$ $\textbf{0.046}\pm 0.004$ $\textbf{0.060}\pm 0.005$ $\textbf{0.071}\pm 0.005$ $\textbf{0.081}\pm 0.006$
($33\pm 0.7$) ($43\pm 0.76$) ($52\pm 0.6$) ($59\pm 0.6$)
(Li et al., 2023) $3.1\cdot 10^{-4}\pm 3.3\cdot 10^{-5}$ $0.081\pm 0.008$ $0.106\pm 0.001$ $0.128\pm 0.012$ $0.150\pm 0.014$
($61\pm 5.9$) ($80\pm 7.4$) ($97\pm 9.2$) ($114\pm 10.8$)
$0.5$ $\text{IC}(0)$ $7.4\cdot 10^{-5}\pm 2.5\cdot 10^{-6}$ $\textbf{0.052}\pm 0.004$ $\textbf{0.066}\pm 0.005$ $\textbf{0.078}\pm 0.006$ $\textbf{0.088}\pm 0.007$
($37\pm 1.2$) ($48\pm 1.1$) ($57\pm 1.2$) ($64\pm 1.2$)
(Li et al., 2023) $4.1\cdot 10^{-4}\pm 1.3\cdot 10^{-5}$ nan1 nan1 nan1 nan1